OpenStudy (anonymous):

Hi. Im doing a worksheet for my final exam and I have unfortunately forgotten how to answer some questions. The questions are as followed for the equation N2 + 3H2 -> 2NH3: a. If 6.100mol of hydrogen were to react, how many moles of nitrogen would be needed? b. If 25.70 grams of nitrogen were to react, how many grams of ammonia would be produced?

OpenStudy (anonymous):

N2+3H2---->2NH3 The molar ratio of N2:H2 is 1:3 this will help u to find the answers for Question a) and b).

OpenStudy (shubhamsrg):

So you have unfortunately forgotten how to answer some questions . LOL ! :D

OpenStudy (anonymous):

a. \[? mol N2 = 6.100 mol H2 \times (\frac{ 1 mol N2 }{ 3mol H2 } )\] b. \[?gr NH3 = 25.70gr N2 \times (\frac{ 1 mol N2 }{28 gr N2}) \times (\frac{ 1mol NH3 }{ 1mol N2 }) \times (\frac{ 17 gr NH3 }{ 1mol NH3 }) \]